# From quantum enhanced to quantum inspired Monte Carlo

### v1

@@ III Optimal parameter regimes
The first objective of the manuscript is to gain a deeper understanding of the parameter regimes in which the quantum proposal strategy is effective. We compute the spectral gap exactly using exact continuous-time limit integration of the quantum dynamics across a wide range of $\gamma$ and $t$ values for varying system sizes. Fig.1 (panel a) displays the landscape of $\delta$ values for a $n=9$ qubit system. A similar grid search for smaller systems reveals consistent results.
In general, the relationship between $\delta$ and the parameters $\gamma$ and $t$ in Fig.1 (panel a) appears complex, but some intuition can be gained by examining the limiting cases, similarly to Ref. [45]. When $\gamma=1$, the Hamiltonian consists solely of the mixing term, leading to a single-qubit rotation around the x-axis:
@@
This produces oscillatory behavior along the $t$-axis. When measured in the computational basis, this quantum evolution works as a classical proposal, flipping each spin with a probability of $\sin^{2}(t)$. For example, at $t=\frac{\pi}{4}$, the process reduces to the random proposal strategy. On the other hand, for sufficiently small $\gamma$, the landscape can be understood using perturbation theory [11]. The evolution under the classical Hamiltonian $H_{c}$ perturbed with a small mixing term $H_{\textrm{mix}}$ generates transitions between computational eigenstates |$\ket{j}$ and |$\ket{k}$ with a probability given by:
$P(\ket{j}\rightarrow\ket{k})\propto|\bra{k}H_{\textrm{mix}}\ket{j}|^{2}\gamma^{2}+\mathcal{O}(\gamma^{3}).$ (14)
@@ III.1 Optimal $\gamma$
Remarkably, there is a pronounced region of high spectral gap for $\gamma$ values in the range $[0.3,0.5]$. Averaging over $t$ in the range $[12,30]$, we observe peaks in $\delta$ across all system sizes, as illustrated in Fig. 1 (panel b). The corresponding optimal $\gamma_{\textrm{opt}}$ values are plotted for each system size in Fig. 1 (panel c). We see that $\gamma_{\textrm{opt}}$ decreases monotonically with system size $n$, flattening out around $\gamma=0.42$ for the $n=9$ system. The apparent convergence to an asymptotic value is physically intuitive, although verification would require larger-scale simulations. If an asymptotically optimal value indeed exists, the quantum proposal strategy could be simplified and enhanced by using a fixed optimal $\gamma$ value rather than picking $\gamma$ at random.
We observe that values of $\gamma$ leading to the highest spectral gaps lie in the region close to the expected phase transition of the quantum spin glass model [41, 46, 42]. To validate this observation, we compute the phase diagram with respect to $\gamma$ and $T$ in Sect. A and determine the critical gamma value $\gamma_{c}=0.50\pm 0.02$. This value is close, but larger than the optimal $\gamma_{\textrm{opt}}$ determined by grid search. Therefore, we conjecture that critical behavior near the phase transition, characterized by enhanced quantum fluctuations, helps to generate a good proposal distribution. However, the exact mechanism underlying this relationship remains unclear and needs further investigation.
@@ III.2 Total evolution time
To understand the role of the total evolution time, we fix $\gamma$ at its optimal value for each size of the system and compute the absolute spectral gap as a function of total evolution time. As shown in Fig. 2 (panel $a$), the gap increases rapidly at short times before stabilizing and fluctuating around a steady value at longer times. This suggests that, beyond a certain point, the gap becomes only weakly time-dependent.
Figure 2: (a) Spectral gap as a function of the total evolution time $t$. Systems of sizes $n=3$ to $n=9$ are shown. (b) Time $t_{\mathrm{0.01}}$ to ”hit” a fixed gap value $\delta=0.01$ as a function of system size $n$.
A key question is how the time to reach this steady value scales with system size. Due to significant fluctuations, precise measurement of the convergence time is challenging. Instead, we calculate the first time the gap crosses a fixed value, $\delta=0.01$, for all system sizes. The results, shown in Fig. 2 (panel $b$), indicate linear scaling, suggesting that the quantum speedup remains unaffected asymptotically (see Sect. II.4)
@@
Empirically, we find that setting $\gamma$ to a fixed value $\gamma=0.45$, slightly below its critical value, and setting fixed $t=12$, slightly outperforms the original fully randomized strategy, as shown in Fig.3.
@@
While these observations cannot be considered conclusive as they are based on a fairly limited number of spins, we do not identify any clear exponential bottleneck that may have been previously overlooked. Overall, it is interesting to notice that the randomized strategy is already competitive with respect to a more fine-tuned parameter choice.
@@ IV.1 Time-dependent Hamiltonian evolution
The surprisingly good performance of the original randomized strategy, challenges us to explore different types of circuits. The quantum proposal method is not limited to time-independent Hamiltonians. Time-dependent Hamiltonians can also be employed for time evolution, provided they satisfy the symmetry constraint $U(\tau)=U(\tau)^{T}$, where $\tau$ represents the total evolution time. Ref. [11] demonstrates that a sufficient condition is $H(t)=H(t)^{T}$ and $H(t)=H(\tau-t)$ for all $t\in[0,\tau]$. In the following, we introduce the time-dependence by allowing $\gamma(t)$ to vary with time:
$H(t)=(1-\gamma(t))\alpha H_{c}+\gamma(t)H_{\textrm{mix}}$ (15)
@@
Therefore, in this study, we adopt an uninformed hands-off approach by using a classical optimizer to design the optimal $\gamma(t)$ schedule, inspired by recent variational quantum adiabatic algorithms for ground state preparation [47, 48, 49, 50]. These algorithms use classical optimization to adjust adiabatic evolution parameters to the Hamiltonian gap structure. For example, by changing the annealing speed to minimize Landau-Zener transitions when the evolution passes through the region with the minimal spectral gap.
However, we are addressing a fundamentally different problem here. Landau-Zener transitions, which are typically detrimental in the context of adiabatic quantum evolution, can be a valuable resource in our case if they occur between states that are close in energy but potentially far in Hamming distance [11]. Thus, it is worth investigating whether a classical optimizer can identify and leverage such diabatic effects by adjusting the time-evolution shape of our Hamiltonian [51]. In this context, we consider employing Bayesian optimization (BO) to design an optimized schedule for $\gamma(t)$. BO is a widely used global optimization technique [52], particularly effective when the cost function is expensive to evaluate, such as in computing the absolute spectral gap. We refer the reader to Appendix B for a brief overview of BO, and more details.
The spectral gaps achieved with these optimized schedules are presented in Fig. 4. BO-optimized schedules outperform the time-independent strategy with fixed $t$ and $\gamma$ across all system sizes. However, the performance gains are irregular, and the resulting scaling coefficient is, in fact, worse. For $n\in[8,9]$, BO-optimized schedules closely mimic the time-independent strategy. A key distinction is that the optimized schedules ramp up the transverse field gradually, whereas the time-independent approach applies it instantaneously.
Figure 4: The absolute spectral gap $\delta$ for the fixed $t=12$ and $\gamma=0.45$ quantum proposal (purple) and the quantum proposal based on the classically optimized time-dependent Hamiltonian (dark grey). Exponential fits are shown as dotted lines with the corresponding fit parameters.
Our findings indicate that the classical optimizer can indeed surpass the original strategy for smaller systems, potentially by leveraging strong diabatic effects. However, for larger systems, the optimizer tends to converge to a similar time evolution as the original strategy. Overall, the fact that this methods perform irregularly, leads to a worse scaling.
@@ IV.2 Symmetric QAOA circuit
An alternative way of constructing the quantum proposal is utilizing a parametrized, symmetric ansatz. Ref. [14] proposes a variational circuit inspired by QAOA, and test it against classical update. However, the extension is not tested against the original QeMC proposal, so we need to understand if this choice actually constitutes and improvement over the original method.
Figure 5: QAOA inspired circuit. Variation of the spectral gap with system size $n$. Panel (a) corresponds to a circuit depth of $p=5$, panel (b) to $p=20$, and panel (c) to $p=50$. An increase in circuit depth from $p=5$ to $p=20$ results in a small improvement in the spectral gap, but no further improvement is observed when increasing the depth to $p=50$. Each plot represents data obtained from 100 random instances. The parameters used for the quantum strategy are $\gamma=0.45$, $t=12$, and $T=1$.
The ansatz has the following form:
$U=V(\boldsymbol{\beta},\boldsymbol{\gamma})^{\top}V(\boldsymbol{\beta},\boldsymbol{\gamma}),$ (16)
@@
$V(\boldsymbol{\beta},\boldsymbol{\gamma})=U_{C}\left(\gamma_{p}\right)U_{B}\left(\beta_{p}\right)\cdots U_{C}\left(\gamma_{1}\right)U_{B}\left(\beta_{1}\right),$ (17)
with $U_{B}(\beta)=e^{-iH_{\textrm{mix}}\beta}$, and $U_{C}(\gamma)=e^{-i\alpha H_{c}\gamma}$. It features $p$ layers, implying that there are $2p$ parameters to optimize: $\beta=\{\beta_{1},\cdots,\beta_{p}\}$ and $\gamma=\{\gamma_{1},\cdots,\gamma_{p}\}$. This circuit resembles the structure of a discrete reverse-annealing process, which has also been proposed in Ref. [11] as a possible quantum process for the method. However, this circuit can be more general and, most importantly, can offer a constant-depth ansatz.
Following Ref. [14], we simplify the optimization process, using the following constraint $\theta=\beta_{1}=\cdots=\beta_{p}=\gamma_{1}=\cdots=\gamma_{p}.$ This is necessary as optimizing from scratch QAOA parameters without any prior guess [53] has proven to be inefficient, especially when the cost function comes with a statistical error bar [54].
Unlike plain QAOA, where the cost function is an energy, the objective here is to minimize the mixing time of the MCMC. The practical approach proposed in Ref. [14] is to perform short MCMC runs and compute the acceptance rate. While this argument is not rigorous, the authors of Ref. [14] found a correlation between delta and the acceptance rate, which holds only for small values of theta. Therefore, the practical strategy would consist of minimizing the acceptance rate (AR) under these conditions. In turn, the AR can be measured empirically from MCMC runs, or exactly from the definition of the $Q$ and $A$ matrices, though this would be exponentially costly.
However, we expect that directly maximizing $\delta$ will lead to the best result the algorithm can offer, as opposed to minimizing the acceptance rate. Since our goal is to assess whether the entire QAOA ansatz is beneficial at all, we also chose to try this second objective. By directly maximizing $\delta$, as in the previous Sections, we aim to test the performance of the variational circuit without further assumptions. While also not scalable, this approach provides a more straightforward benchmark for the small system sizes considered here. In this Section, we perform statevector simulation of the circuits using Qiskit software package [55]. We plot the results in Fig.5, showcasing the performance for various depths $p$. While there is a slight improvement in the gap between $p=5$ and $p=20$, it is evident that, in this case as well, the scaling of the gap versus system size is not improved compared to the original approach.
@@
In the previous Sections, we observed the remarkable, empirical superiority of the original Hamiltonian dynamics approach. This Section explores whether this protocol is robust against algorithmic and approximation errors. The first point allows us to determine whether we truly need to perform the quantum dynamics simulation in its continuous-time limit, thus saving resources. The second point could establish a new class of classical update schemes that are only inspired by the QeMC protocol but do not require quantum hardware.
@@ V.1 Trotter error
Figure 6: (a) Impact of trotter step size $dt$ on the quantum proposal strategy. (b) Objective function to find optimal trotter step size $dt_{opt}$ for different system sizes. The optimal step size converges to $dt=0.8$ for $n=9$. Here the quantum and quantum trotter strategy use $\gamma=0.45$ and $t=12$.
@@
In the original paper, a Trotter time step of 0.8 was chosen, primarily to meet hardware constraints  [11]. Indeed, it has been shown, both theoretically [56] and experimentally  [11, 57], that a finite time step is optimal in the presence of hardware gate noise. Moreover, a large time step allowed for the practical implementation of the $e^{-i\theta/2\leavevmode\nobreak\ Z\otimes Z}$ gate, using pulse-efficient cross-resonance gates  [58].
Here, we run the QeMC algorithm using $t=12$ and $\gamma=0.45$, using discrete-time evolution methods. We vary the Trotter step $dt$ and analyze the gap scaling. The results are reported in Fig.  6 (panel a). We observe that different values of $dt$ lead to a spectrum of results. Large $dt$’s yield a scaling similar to Uniform, inefficient proposals, while we recover the continuous-time result only when $dt\rightarrow 0$. Interestingly, finite $dt$ never produces better scaling than the continuous-time result.
There is, however, a subtlety here. While smaller $dt$’s improve the scaling, they require more Trotter steps, $m$, to reach a target $t=m\leavevmode\nobreak\ dt$. This remains true even when $t$ is sampled randomly. To identify the computationally optimal $dt$, we maximize the following objective function:
$f(dt;n)=\frac{\delta(dt;n)\leavevmode\nobreak\ dt}{t},$ (18)
where $t=12$ is a constant. Eq. 18 follows from the definition of $\delta$ as the inverse mixing time, and that the computational cost grows with decreasing $dt$. We observe that this objective function shows indeed a maximum, whose position depends on $n$. Results in Fig. 6(panel b) suggest that the optimal time steps concentrate, for the larger $n$’s, around the value of $dt_{opt}=0.8$. Surprisingly, this is the same value chosen for the hardware simulation of Ref. [11], though for different motivations. We receive the same value for $dt_{opt}$ when running the QeMC algorithm in its original setting, i.e., randomizing over the total evolution time $t$, and $\gamma$.

### v2

@@ III Optimal parameter regimes
The first objective of the paper is to gain a deeper understanding of the parameter regimes in which the quantum proposal strategy is effective. We perform the analysis without actual quantum hardware by computing the spectral gap numerically using exact continuous-time limit integration of the quantum dynamics across a wide range of $\gamma$ and $t$ values for varying system sizes. Fig.1 (panel a) displays the landscape of $\delta$ values for a $n=9$ qubit system. A similar grid search for smaller systems reveals consistent results.
In general, the relationship between $\delta$ and the parameters $\gamma$ and $t$ in Figure1 (a) appears complex, but some intuition can be gained by examining the limiting cases, similarly to Ref. [44]. When $\gamma=1$, the Hamiltonian consists solely of the mixing term, leading to a single-qubit rotation around the x axis:
@@
This produces oscillatory behavior along the $t$ axis. When measured in the computational basis, this quantum evolution works as a classical proposal, flipping each spin with a probability of $\sin^{2}(t)$. For example, at $t=\frac{\pi}{4}$, the process reduces to the random proposal strategy. On the other hand, for sufficiently small $\gamma$, the landscape can be understood using perturbation theory [11]. The evolution under the classical Hamiltonian $H_{c}$ perturbed with a small mixing term $H_{\textrm{mix}}$ generates transitions between computational eigenstates |$\ket{j}$ and |$\ket{k}$ with a probability given by:
$P(\ket{j}\rightarrow\ket{k})\propto|\bra{k}H_{\textrm{mix}}\ket{j}|^{2}\gamma^{2}+O(\gamma^{3}).$ (14)
@@ III.1 Optimal $\gamma$
Remarkably, there is a pronounced region of high spectral gap for $\gamma$ values in the range $[0.3,0.5]$. Averaging over $t$ in the range $[12,30]$, we observe peaks in $\delta$ across all system sizes, as illustrated in Fig. 1 (b). The corresponding optimal $\gamma_{\textrm{opt}}$ values are plotted for each system size in Fig. 1 (c). We see that $\gamma_{\textrm{opt}}$ decreases monotonically with system size $n$, flattening out around $\gamma=0.42$ for the $n=9$ system. The apparent convergence to an asymptotic value is physically intuitive, although verification would require larger-scale simulations. If an asymptotically optimal value indeed exists, the quantum proposal strategy could be simplified and enhanced by using a fixed optimal $\gamma$ value rather than picking $\gamma$ at random.
We observe that values of $\gamma$ leading to the highest spectral gaps lie in the region close to the expected phase transition of the quantum-spin-glass model [40, 45, 41]. To validate this observation, we compute the phase diagram with respect to $\gamma$ and $T$ in Appendix A and determine the critical gamma value $\gamma_{c}=0.50\pm 0.02$. This value is close to but larger than the optimal $\gamma_{\textrm{opt}}$ determined by grid search. Therefore, we conjecture that critical behavior near the phase transition, characterized by enhanced quantum fluctuations, helps to generate a good proposal distribution. However, the exact mechanism underlying this relationship remains unclear and needs further investigation.
@@ III.2 Total evolution time
To understand the role of the total evolution time, we fix $\gamma$ at its optimal value for each size of the system and compute the absolute spectral gap as a function of total evolution time. As shown in Fig. 2 (a), the gap increases rapidly at short times before stabilizing and fluctuating around a steady value at longer times. This suggests that, beyond a certain point, the gap becomes only weakly time dependent.
Figure 2: (a) Spectral gap as a function of the total evolution time $t$. Systems of sizes $n=3-9$ are shown. (b) Time $t_{\mathrm{0.01}}$ to hit a fixed gap value $\delta=0.01$ as a function of system size $n$.
A key question is how the time to reach this steady value scales with system size. Due to significant fluctuations, precise measurement of the convergence time is challenging. Instead, we calculate the first time the gap crosses a fixed value, $\delta=0.01$, for all system sizes. The results, shown in Fig. 2 (b), indicate linear scaling, suggesting that the quantum speedup remains unaffected asymptotically (see Sec. II.4)
@@
Empirically, we find that setting $\gamma$ to a fixed value $\gamma=0.45$, slightly below its critical value, and setting fixed at $t=12$ slightly outperforms the original fully randomized strategy, as shown in Fig. 3.
@@
While these observations cannot be considered conclusive as they are based on a fairly limited number of spins, we do not identify any clear exponential bottleneck that may have been previously overlooked. Overall, it is interesting to note that the randomized strategy is already competitive with respect to a more fine-tuned parameter choice.
@@ IV.1 Time-dependent Hamiltonian evolution
The surprisingly good performance of the original randomized strategy, challenges us to explore different types of circuits. The quantum proposal method is not limited to time-independent Hamiltonians. Time-dependent Hamiltonians can also be employed for time evolution, provided they satisfy the symmetry constraint $U(\tau)=U(\tau)^{T}$, where $\tau$ represents the total evolution time. Reference  [11] demonstrates that a sufficient condition is $H(t)=H(t)^{T}$ and $H(t)=H(\tau-t)$ for all $t\in[0,\tau]$. In the following, we introduce the time-dependence by allowing $\gamma(t)$ to vary with time:
$H(t)=[1-\gamma(t)]\alpha H_{c}+\gamma(t)H_{\textrm{mix}}$ (15)
@@
Therefore, in this study, we adopt an uninformed hands-off approach by using a classical optimizer to design the optimal $\gamma(t)$ schedule, inspired by recent variational quantum adiabatic algorithms for ground state preparation [46, 47, 48, 49]. These algorithms use classical optimization to adjust adiabatic evolution parameters to the Hamiltonian gap structure, for example, by changing the annealing speed to minimize Landau-Zener transitions when the evolution passes through the region with the minimal spectral gap.
However, we are addressing a fundamentally different problem here. Landau-Zener transitions, which are typically detrimental in the context of adiabatic quantum evolution, can be a valuable resource in our case if they occur between states that are close in energy but potentially far in Hamming distance [11]. Thus, it is worth investigating whether a classical optimizer can identify and leverage such diabatic effects by adjusting the time-evolution shape of our Hamiltonian [50]. In this context, we consider employing Bayesian optimization (BO) to design an optimized schedule for $\gamma(t)$. Bayesian optimization is a widely used global optimization technique [51], particularly effective when the cost function is expensive to evaluate, such as in computing the absolute spectral gap. We refer the reader to Appendix B for a brief overview of BO and more details.
The spectral gaps achieved with these optimized schedules are presented in Fig. 4. Bayesian-optimization-optimized schedules outperform the time-independent strategy with fixed $t$ and $\gamma$ across all system sizes. However, the performance gains are irregular and the resulting scaling coefficient is in fact worse. For $n\in[8,9]$, BO-optimized schedules closely mimic the time-independent strategy. A key distinction is that the optimized schedules ramp up the transverse field gradually, whereas the time-independent approach applies it instantaneously.
Figure 4: Absolute spectral gap $\delta$ for the fixed $t=12$ and $\gamma=0.45$ quantum proposal (purple) and the quantum proposal based on the classically optimized time-dependent Hamiltonian (dark grey). Exponential fits are shown as dotted lines with the corresponding fit parameters.
Our findings indicate that the classical optimizer can indeed surpass the original strategy for smaller systems, potentially by leveraging strong diabatic effects. However, for larger systems, the optimizer tends to converge to a time evolution similar to that of the original strategy. Overall, the fact that this method performs irregularly leads to worse scaling.
@@ IV.2 Symmetric QAOA circuit
An alternative way of constructing the quantum proposal is utilizing a parametrized, symmetric ansatz. Reference  [14] proposes a variational circuit inspired by the quantum approximate optimization algorithm (QAOA) and tests it against classical update. However, the extension is not tested against the original QEMC proposal, so we need to understand if this choice actually constitutes an improvement over the original method.
Figure 5: QAOA-inspired circuit. The variation of the spectral gap with system size $n$ is plotted for circuit depths (a) $p=5$, (b) $p=20$ and (c) $p=50$. An increase in circuit depth from $p=5$ to $20$ results in a small improvement in the spectral gap, but no further improvement is observed when increasing the depth to $p=50$. Each plot represents data obtained from 100 random instances. The parameters used for the quantum strategy are $\gamma=0.45$, $t=12$, and $T=1$.
The ansatz has the form
$U=V(\bm{\beta},\bm{\gamma})^{\top}V(\bm{\beta},\bm{\gamma}),$ (16)
@@
$V(\bm{\beta},\bm{\gamma})=U_{C}\left(\gamma_{p}\right)U_{B}\left(\beta_{p}\right)\cdots U_{C}\left(\gamma_{1}\right)U_{B}\left(\beta_{1}\right),$ (17)
with $U_{B}(\beta)=e^{-iH_{\textrm{mix}}\beta}$ and $U_{C}(\gamma)=e^{-i\alpha H_{c}\gamma}$. It features $p$ layers, implying that there are $2p$ parameters to optimize: $\beta=\{\beta_{1},\cdots,\beta_{p}\}$ and $\gamma=\{\gamma_{1},\cdots,\gamma_{p}\}$. This circuit resembles the structure of a discrete reverse-annealing process, which has also been proposed in Reference  [11] as a possible quantum process for the method. However, this circuit can be more general and, most importantly, can offer a constant-depth ansatz.
Following Ref. [14], we simplify the optimization process, using the constraint $\theta=\beta_{1}=\cdots=\beta_{p}=\gamma_{1}=\cdots=\gamma_{p}.$. This is necessary as optimizing from scratch QAOA parameters without any prior guess [52] has proven to be inefficient, especially when the cost function comes with a statistical error bar [53].
Unlike the plain QAOA, where the cost function is an energy, the objective here is to minimize the mixing time of the MCMC. The practical approach proposed in Ref. [14] is to perform short MCMC runs and compute the acceptance rate. While this argument is not rigorous, the authors of Ref. [14] found a correlation between $\delta$ and the acceptance rate, which holds only for small values of $\theta$. Therefore, the practical strategy would consist of minimizing the acceptance rate (AR) under these conditions. In turn, the AR can be measured empirically from MCMC runs, or exactly from the definition of the $Q$ and $A$ matrices, though this would be exponentially costly.
However, we expect that directly maximizing $\delta$ will lead to the best result the algorithm can offer, as opposed to minimizing the acceptance rate. Since our goal is to assess whether the entire QAOA ansatz is beneficial at all, we also choose to try this second objective. By directly maximizing $\delta$, as in the previous sections, we aim to test the performance of the variational circuit without further assumptions. While also not scalable, this approach provides a more straightforward benchmark for the small system sizes considered here. In this section, we perform Statevector simulation of the circuits using qiskit software package [54]. We plot the results in Fig.5, showcasing the performance for various depths $p$. While there is a slight improvement in the gap between $p=5$ and $20$, it is evident that, in this case as well, the scaling of the gap versus system size is not improved compared to the original approach.
@@
In the previous sections, we observed the remarkable empirical superiority of the original Hamiltonian dynamics approach. This section explores whether this protocol is robust against algorithmic and approximation errors. The first point allows us to determine whether we truly need to perform the quantum dynamics simulation in its continuous-time limit, thus saving resources. The second point could establish a new class of classical update schemes that are only inspired by the QEMC protocol but do not require quantum hardware.
@@ V.1 Trotter error
Figure 6: (a) Impact of Trotter step size $dt$ on the quantum proposal strategy. (b) Objective function to find optimal Trotter step size $dt_{\text{opt}}$ for different system sizes. The optimal step size converges to $dt=0.8$ for $n=9$. Here the quantum and quantum Trotter strategies use $\gamma=0.45$ and $t=12$.
@@
In the original paper, a Trotter time step of 0.8 was chosen, primarily to meet hardware constraints  [11]. Indeed, it has been shown, both theoretically [55] and experimentally  [11, 56], that a finite time step is optimal in the presence of hardware gate noise. Moreover, a large time step allowed for the practical implementation of the $e^{-i\theta/2~{}Z\otimes Z}$ gate, using pulse-efficient cross-resonance gates  [57].
Here we run the QEMC algorithm using $t=12$ and $\gamma=0.45$, using discrete-time evolution methods. We vary the Trotter step $dt$ and analyze the gap scaling. The results are reported in Fig. 6 (a). We observe that different values of $dt$ lead to a spectrum of results. Large values of $dt$ yield a scaling similar to uniform inefficient proposals, while we recover the continuous-time result only when $dt\rightarrow 0$. Interestingly, finite $dt$ never produces better scaling than the continuous-time result.
There is however a subtlety here. While smaller values of $dt$ improve the scaling, they require more Trotter steps $m$ to reach a target $t=m~{}dt$. This remains true even when $t$ is sampled randomly. To identify the computationally optimal $dt$, we maximize the objective function
$f(dt;n)=\frac{\delta(dt;n)~{}dt}{t},$ (18)
where $t=12$ is a constant. Equation 18 follows from the definition of $\delta$ as the inverse mixing time and the computational cost growing with decreasing $dt$. We observe that this objective function shows indeed a maximum, whose position depends on $n$. Results in Fig. 6(b) suggest that the optimal time steps concentrate, for the larger $n$, around the value of $dt_{\text{opt}}=0.8$. Surprisingly, this is the same value chosen for the hardware simulation of Ref. [11], though for different motivations. We receive the same value for $dt_{\text{opt}}$ when running the QEMC algorithm in its original setting, i.e., randomizing over the total evolution time $t$ and $\gamma$. The simulations in these sections were done using the Julia programming language [58] and using 48 CPUs with 2560 MB memory each of the ETH Euler Cluster. Individual instances were calculated in parallel using multi-processing and the Distributed.jl library .